OpenStudy (anonymous):

Confused after distributing. 4(-5y+1)= -3(-y +6). If you can show the steps after distributing would help me a lot.

OpenStudy (anonymous):

After you distribute, solve for y

OpenStudy (anonymous):

-20y + 4 = 3y- 6 Get the Variables on one side and numbers on the other side -20y + 20y + 4 = 3y - 6 + 20y Simplify 4 = 23y - 6 4+6 = 23y -6+6 10=23y 10/23 = 23y/23 10/23=y Chech Your answer by substituting 10/23 into y iin the equation

OpenStudy (anonymous):

Thank you! ;]